OpenStudy (anonymous):

Express the required measurement in exact form. What is the degree measure of an arc 4pi ft. long in a circle of radius 10 ft.? help me please

OpenStudy (anonymous):

arc length =4pi circle circumference =20pi {circumference of circle=2pi times radius} therefore 1/5 of the circle circumference is covered by the arc . therefore 360/5 =72 deg.

OpenStudy (anonymous):

sure. way two radian measure of angle is arc length divided by radius. in your case \[\frac{4\pi}{10}=\frac{2\pi}{5}\] convert to degrees via \[\frac{180}{\pi}\times \frac{2\pi}{5}=72\]